OpenStudy (anonymous):

y-3x=13 I don't know how to do this :(

OpenStudy (triciaal):

what are you suppose to do ? make y the subject? we really can't help if you don't put the instructions.

OpenStudy (anonymous):

Solve the literal equation for y

OpenStudy (anonymous):

It just means have y alone equal to an expression. y=mx+n

OpenStudy (anonymous):

So all you need to do in the equation you posted is add 3x to both sides.

OpenStudy (anonymous):

which would be 9x and 21?

OpenStudy (anonymous):

y-3x=13 y-3x+3x=13+3x y+0=13+3x y=3x+13

OpenStudy (anonymous):

Sorry I don't understand how to do this at all :/

OpenStudy (anonymous):

Do you know how to solve basic equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Can you solve this for x? 2x=4

OpenStudy (anonymous):

4-2

OpenStudy (anonymous):

= 2x

OpenStudy (anonymous):

To have x alone you have to divide both sides by 2

OpenStudy (anonymous):

2x divided by 2 is x 4 divided by 2 is 2 2x=4 /:2 x=2

OpenStudy (anonymous):

Oh thanks :)

OpenStudy (anonymous):

Now do this for me: 3x-1=5

OpenStudy (anonymous):

Solve for x, make sure it's alone at the end.

OpenStudy (anonymous):

3x+3x= 5+3x 0x+ 1 = 5+3x 5+3x=8+1+ 9x ?

OpenStudy (anonymous):

3x-1=5 / add 1 to both sides 3x-1+1=5+1 3x+0=6 3x=6 /divide both sides by 3 to have x alone x=2

OpenStudy (anonymous):

Remember that you can't add 2x+3 to have 6x, they are different factors

OpenStudy (anonymous):

I don't get it still

OpenStudy (anonymous):

On one side you want numbers with x attached to them, on the other just numbers

OpenStudy (anonymous):

Do this expression, not an equation this time. Just add those up. 2x+4x+5+7

OpenStudy (anonymous):

18

OpenStudy (anonymous):

You can only add up numbers with an x, and regular numbers

OpenStudy (anonymous):

6x+12

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Do another expression: 5x+2-3x+1

OpenStudy (anonymous):

8x+3

OpenStudy (anonymous):

Notice the minus on the 3x

OpenStudy (anonymous):

Algebra 1 is too hard:/

OpenStudy (anonymous):

If you want, you can rewrite them, the order doesn't matter in an expression

OpenStudy (anonymous):

5x-3x+2+1 2x+3

OpenStudy (anonymous):

It's easier to see it that way.

OpenStudy (anonymous):

2y-18x=-26

OpenStudy (anonymous):

Add 18x to both sides

OpenStudy (anonymous):

18x=-26 +18x =18x 0x 34x

OpenStudy (anonymous):

then divide 2 to both sides?

OpenStudy (anonymous):

its wrong isn't it

OpenStudy (anonymous):

2y-18x+18x=-26+18x 2y+0=26+18x 2y=26+18x now divide by 2, yes.

OpenStudy (anonymous):

52 and 17?

OpenStudy (anonymous):

If you divide every individual bit by 2, what do you get 2y, 26 and 18x

OpenStudy (anonymous):

2y=26+18x /:2 y=13+6x does that make sense?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

Which is why i wanted to practice some more basic things before we move to the exercises you show me.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In equations, the point is to get the numbers with the x to one side, and regular numbers to the other side.

OpenStudy (anonymous):

4x+2=1+3x

OpenStudy (anonymous):

To move them around, we need to either add or subtract that same number

OpenStudy (anonymous):

For example, to move that 3x with the 4x

OpenStudy (anonymous):

We need to subtract 3x from both sides

OpenStudy (anonymous):

- 2 to both sides?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

4x+2-3x=1+3x-3x

OpenStudy (anonymous):

+3x-3x is equal to zero, so it gets cancelled out

OpenStudy (anonymous):

4x+2-3x=1

OpenStudy (anonymous):

Now we need to move the two, move it for me.

OpenStudy (anonymous):

2 -2

OpenStudy (anonymous):

4x+2-3x=1 subtract 2 from both sides of the equation

OpenStudy (anonymous):

simply add -2 to both sides

OpenStudy (anonymous):

2 =3x -2 -2

OpenStudy (anonymous):

4x+2-2-3x=1-2 4x-3x=-1 1x=-1

OpenStudy (anonymous):

Also, writing x and 1x is the same.

OpenStudy (anonymous):

Let's try another one.

OpenStudy (anonymous):

3x+1=2 How would you have 3x alone in here?

OpenStudy (anonymous):

-1 to both sides 3x+1=2 3x -1= -1 3x= 0=1 x=4

OpenStudy (anonymous):

Remember that you are removing 1 from both sides without changing any numbers, just tuck them in there

OpenStudy (anonymous):

3x+1-1=2-1 3x=1

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

Now if you wanted x alone without the 3, you would have to divide both sides by 3

OpenStudy (anonymous):

3x divided by 3 is x 1 divided by 3 is a third

OpenStudy (anonymous):

\[3x=1\] \[x=\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

I think i'm starting to get it now

OpenStudy (anonymous):

Great :)

OpenStudy (anonymous):

5x-4=10-2x

OpenStudy (anonymous):

5x -4=10-2x 5x +4= +4-2x 5x +0= 14-2x 5x=0=12

OpenStudy (anonymous):

I get the beginning of it I think but idk the ending

OpenStudy (anonymous):

5x-4=10-2x 5x-4+4=10-2x+4 you added 4 right?

OpenStudy (anonymous):

I did the first time and got 8x

OpenStudy (anonymous):

10-8x

OpenStudy (anonymous):

There is no need to show that -4+4 is 0, just eliminate it from the start Reorder the positions so you see what's numbers and what's x's 5x=10+4-2x

OpenStudy (anonymous):

5x=14-2x

OpenStudy (anonymous):

5x=12x

OpenStudy (anonymous):

14 is a number, 2x is an x

OpenStudy (anonymous):

we need to move that -2x to the x's side.

OpenStudy (anonymous):

In other words, to have 14 alone, we need to find a way how can we have a zero there

OpenStudy (anonymous):

If we add 2x it's going to be 14-2x+2x 14-0 14

OpenStudy (anonymous):

But since we added 2x, we must also add it on the other side

OpenStudy (anonymous):

5x+2x=14-2x+2x 7x=14-0 7x=14

OpenStudy (anonymous):

How would you isolate x in here?

OpenStudy (anonymous):

divide it by 7x

OpenStudy (anonymous):

If you were to divide by 7x, you would get this: \[\frac{ 7x }{ 7x }=\frac{ 14 }{ 7x }\] \[1 = \frac{ 2 }{ x }\]

OpenStudy (anonymous):

But since we still want x, we only divide by 7.

OpenStudy (anonymous):

\[\frac{ 7x }{ 7 } = \frac{ 14 }{ 7 }\] x=2

OpenStudy (anonymous):

Isolate x in here: 3x=9

OpenStudy (anonymous):

3

OpenStudy (anonymous):

x is indeed equal to 3

OpenStudy (anonymous):

2y+3y+8x-3x+1+4